# Pareto-Optimal Learning from Preferences with Hidden Context

### v1

@@ Abstract
Ensuring AI models align with human values is essential for their safety and functionality. Reinforcement learning from human feedback (RLHF) uses human preferences to achieve this alignment. However, preferences sourced from diverse populations can result in point estimates of human values that may be sub-optimal or unfair to specific groups. We propose Pareto Optimal Preference Learning (POPL), which frames discrepant group preferences as objectives with potential trade-offs, aiming for policies that are Pareto-optimal on the preference dataset. POPL utilizes Lexicase selection, an iterative process to select diverse and Pareto-optimal solutions. Our empirical evaluations demonstrate that POPL surpasses baseline methods in learning sets of reward functions, effectively catering to distinct groups without access to group numbers or membership labels. Furthermore, we illustrate that POPL can serve as a foundation for techniques optimizing specific notions of group fairness, ensuring inclusive and equitable AI model alignment.
@@ 1 Introduction
For both safety and functionality, it is critical for AI models to be aligned with the values of human users and stakeholders. Reinforcement learning from human feedback (RLHF) has emerged as an effective mechanism for model alignment, using preferences as a means of capturing human values. However, when preferences are sourced from large sets of potentially diverse people, methods that rely on point estimates of human values are bound to either be sub-optimal with respect to all groups or unfair to certain groups, both problematic in their own ways.
@@
In this work, we build upon the notion of hidden context proposed by Siththaranjan et al. [50] and focus on the problem of Reinforcement Learning from Human Feedback with Hidden Context (RLHF-HC). Hidden context is information that is unavailable to a preference learning system that affects the preferences given. For example, a person’s dominant hand might determine on what side a robotic assistant should hand them an object. Under this formulation, the goal is to build a set of policies that contains the policy that is optimal under the reward function of each group of people with the same hidden context. In practice, we see two clear use cases of such a set of policies. 1) These policies can be selected from at test time to find an optimal policy for a given user without needing to train a new policy and 2) this set can be used to ensure fairness between groups. Minimizing risk with respect to this diverse distribution of policies ensures that no specific group is disregarded in the risk measurement.
@@
Because preferences generated from individuals in different hidden context groups could be contradictory (i.e. not mutually satisfiable by a well-regularized policy or reward function), we propose to frame these preferences as objectives with potential trade-offs between each other. With this re-framing, the optimal policy for each individual hidden context group would be Pareto-optimal (non-dominated) on the dataset of preferences. With this in mind, we propose Pareto Optimal Preference Learning (POPL), where we learn a set of reward functions or policies (directly) that are optimized towards being Pareto-optimal with respect to the set of preferences given by a potentially diverse set of human annotators. To do this, we use an iterative selection process known as Lexicase selection [51], which has been shown under mild assumptions to select individuals that are both Pareto-optimal and diverse. An outline of our method can be found in Figure 1.
We conduct a variety of empirical investigations to show that POPL outperforms strong baselines in learning sets of reward functions. Namely, we show in both a synthetic and large language model (LLM) domain, that POPL is able to cater reward functions for specific groups as well as limit reward hacking and jailbreaking. At the same time, we also show that this technique works for directly catering policies for distinct groups of people, without having access to the number of groups nor membership labels for annotations. Finally, we show that our technique can be used as a precursor for techniques that optimize for particular notions of group fairness.
@@
Figure 1: Outline of our proposed method. Given a set of pairwise preferences over trajectory segments from groups with potentially different ground truth reward functions, we infer a set of reward functions or policies that captures each group’s ground truth, without group membership labels. To do this, we frame reward inference as multi-objective optimization, where each preference forms a single objective, and find a set of Pareto-optimal reward functions or policies.
@@ 2 Related Work
Data used for RLHF systems often comes from multiple people, who are diverse in their preferences and values [7, 45, 5, 49]. This data, when considered in its aggregated form, can not be captured perfectly by a decision-making model that relies on a point estimate of utility [14]. These models try to find a single reward function that is most likely, which is often not the optimal reward function for any one single person. When the groups are not perfectly balanced, the minority groups might be underrepresented in the inferred reward function [50, 24, 37, 43] or simply treated as noise [4]. There have been attempts at explicitly modeling different people with different levels of expertise [29, 19, 30, 3], but these methods generally rely on concrete ways to distinguish between groups.
@@
In the context of RL, Myers et al. [43] outlines an approach to learning a multi-modal reward function from online interaction between a human expert and a preference learning system. Ramé et al. [48] learn a set of reward models by optimizing for diversity amongst the outputs. Whilst similar to our approach, we also aim to align our reward models with hidden context groups through optimizing for Pareto-optimality. For generative AI models and LLMs, there have been a variety of studies attempting to align large models with diverse human preferences. Chakraborty et al. [15] and Siththaranjan et al. [50] learn a mixture of preference distributions or a parameterized reward distribution, respectively. However, both these techniques operate under a contextual bandit setting which results in sub-optimal performance when used in the more general RL setting (discussed further in Section 4). Bradley et al. [9] and Ding et al. [23] leverage fine-tuning to improve the diversity of model responses for better alignment and creativity, which do not directly address the ambiguity and hidden context in human preferences. Jang et al. [34] and Dai et al. [18] elicit preferences specifically along different dimensions in order to cater custom reward functions for users in test time, and to be safe with respect to conflicting objectives, respectively. Whilst we also aim to cater reward functions in test time as well as optimize fairness between groups, we do not have access to labels regarding the context of the preferences generated. Finally, Rame et al. [47] also generates a set of Pareto-optimal reward functions. However, in their setting, the system has access to ground truth reward functions for each group, and the Pareto-front is generated through weight interpolation between these functions.
@@
Bayesian Reward Extrapolation (B-REx) [12] instead learns a distribution of reward models from pairwise human preferences. B-REx is then able to perform Bayesian inference using MCMC [41] to sample from the posterior of reward functions. With this distribution, a practitioner can establish high confidence performance bounds that can be used to assess risk in evaluated policies as well as detect reward hacking behaviors. However, B-REx and other reward inference methods often rely on a faulty assumption that humans provide preferences in a Boltzmann-rational way.
On a separate front, there have been recent efforts on direct, reward-free methods for RLHF [46, 31], where a policy is learned directly from the preferences. However, these methods still implicitly rely on single-point estimates of reward, making them not directly applicable for reward function curation or risk minimization in safety-critical applications with hidden context. Furthermore, the hidden context in these preferences can lead to optimization issues when using these systems [15].
@@ 3 Preliminaries
Reinforcement learning from human feedback considers human preferences over trajectories (or more generally, outputs of a model) in order to learn a reward model or policy that respects the preferences [13, 46, 31, 14, 25].
In order to learn meaningfully from human preferences, one must characterize how preferences are generated from some parameterized preferences model $P(\sigma_{i}\prec\sigma_{j})$. Usually, this preference model is based on the notion of Boltzmann-rationality, where humans generate preferences in accordance to the Bradley-Terry (BT) model [10]. The probability of pairwise preference $(\sigma_{i}\succ\sigma_{j})$ between two trajectories given some utility function $f(\sigma)$ can be written as
@@
where $\beta$ models the confidence in the preference labels. $\beta\rightarrow\infty$ signals that the preference provider is perfectly rational, and $\beta=0$ signals that preferences are random. The Bradley-Terry model is used in many fields, such as psychology [2, 28, 27], However, this model does not perfectly capture the mechanisms driving these preferences [26, 35, 38, 6, 40].
@@
Contrastive Preference Learning (CPL) [31] learns a policy directly from preferences without needing to learn an intermediate reward function. This method uses a regret-based model of preference rather than the standard partial return interpretation. The probability of a preference under a candidate policy can be written as the ratio of the exponentiated sum of log-likelihoods of the chosen segment to the disregarded segment.
Hidden Context
Siththaranjan et al. [50] introduce the problem of preference learning with hidden context. This is the idea that preferences are not only generated based on the exponential utility (partial return or regret), but also on some latent hidden context variable $z$. This variable is not accessible to preference learning systems and poses a challenge as it is often the case that this variable results in breaking the assumption that preferences are generated Boltzmann-rationally.
@@ 4 Problem Statement and Theoretical Foundation
In order to account for hidden context in the preferences learned, [50] introduce Distributional Preference Learning, which relies on a single model of utility $u(s,z)$ to output a distribution of utility assignments $u(z,\cdot)$ for each state $s\in\mathcal{S}$, marginalizing over the hidden context variable $z$. Herein, we will refer to a model that outputs a distribution per state as a Marginalized Distributional Preference Learning (MDPL) system.
However, due to the marginalization process inherent in these systems, the utility function has no concept of persistent annotator identity—the fact that hidden context transcends a single preference annotation. Concretely, assume that users that give the preference $\sigma_{A}\succ\sigma_{B}$, are likely to also give the preference $\sigma_{B}\succ\sigma_{C}$. Being aware of persistent annotator identity would allow for statements to be made like: “the same people that prefer $\sigma_{A}$ to $\sigma_{B}$ are also likely to prefer $\sigma_{B}$ to $\sigma_{C}$.” Mathematically, MDPLs represent the marginalized probability $P(R|s)$ of a specific reward value $R$ in a state $s$. However, to be sensitive to persistent annotator identity, a preference learning framework would need to maintain the full conditional $P(R|s,z)$. In a contextual bandit setting such as those often found for finetuning LLMs [46], this is not an issue, as determining that an output has high risk simply depends on the distribution of rewards attributed to that specific state. However, when learning a policy using this reward function, none of the inter-segment preference dependence is captured. Below, we outline an example to help clarify why this is an issue.
Figure 2: An example of a situation where using POPL is preferable to using a Marginalized Distributional Preference Learning (MDPL) system. Due to the fact that these systems marginalize over the hidden context $z$ for each state, MDPLs are unable to be sensitive to persistent annotator identity. MDPLs represent the distribution of utility values in a column-wise fashion, or maintain a distribution of utilities for each state, that is decoupled from that for other states. POPL, on the other hand, represents the distribution row-wise, finding a set of utility functions that should include the ground truth for each group.
Example 1: MDPLs fail to distinguish between trajectories with different fairness profiles.
Let us consider two groups of people that have differing utilities for states $A,B,C$, as outlined in Figure 2. MDPL systems represent these utilities column-wise, as they maintain a distribution $u(s)$ that has been marginalized over $z$ for every state $s$. Now, consider the three trajectories $AB$, $BC$ and $AC$. These trajectories have different levels of utility when considering individual groups. For example, choosing $AB$ over either of the other two trajectories would have low utility and potentially be unfair (under certain notions of fairness) to an individual from group 2. MDPL systems would not be able to make this distinction, however, ignoring the group-specific utility. Despite the fact that group labels are not available to the preference learning framework, it would be better to maintain the full (un-marginalized) distribution $u(s,z)$ in order to be sensitive to persistent annotator identity. However, as we do not have group labels, we do not have information regarding the number of groups nor the $z$ values that correspond to them. Instead, we can find utility models that satisfy a diverse set of mutually compatible preferences. With enough of these models in a set, we assume that each hidden context group will have its utilities mapped by at least one of the models in the set.
@@
In light of this discussion, we re-frame the problem of preference learning with hidden context as follows. The goal is to learn a set $\Pi=\{\pi_{1},\pi_{2}\cdots,\pi_{n}\}$ of policies such that, for the hidden context group represented by a variable $z\in\mathcal{Z}$, there is a policy $\pi_{z}\in\Pi$ that is the optimal policy for the ground truth reward function $r_{z}$ for the group. Note that this can be accomplished by standard (reward-based) RLHF or direct (reward-free) RLHF. For the standard approach, a series of reward functions $R=\{r_{1},r_{2}\cdots,r_{n}\}$ are first learned from preferences, then used to train $n$ optimal policies. In the direct approach, the policies are learned directly from preferences such as done by Hejna et al. [31] and Rafailov et al. [46].
5 Pareto-Optimality with Lexicase Selection
In this section, we outline an algorithm that aims to generate a series of reward functions (or optimal policies under these reward functions) that align with the preferences of different groups of people. The intuition we rely on for this algorithm is that, in a setting with low noise and sufficient regularization, the set of policies or reward functions that coincide with the various ground truth reward functions (one for each hidden context group) exist on the Pareto front, where each axis corresponds to a single preference.
5.1 Pareto Optimality
@@
Definition 1 (Policy passing preference). A policy $\pi(a|s):\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ passes a preference $(\sigma_{i}\prec\sigma_{j})$ if the probability of the preferred segment $\prod_{(s,a)\in\sigma_{j}}\pi(a|s)$ is higher than the probability of the other segment $\prod_{(s,a)\in\sigma_{i}}\pi(a|s)$. Or, equivalently, if $\sum_{(s,a)\in\sigma_{j}}\log\pi(s,a)>\sum_{(s,a)\in\sigma_{i}}\log\pi(s,a)$.
@@
Definition 2 (Policy-set-relative Pareto-optimality). A policy $\pi(a|s):\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ is Pareto optimal with respect to a set of preferences $\mathcal{P}$ relative to a set of other policies $\Pi=\{\pi_{1},\pi_{2},\cdots,\pi_{n}\}$ if and only if there exists a preference $(\sigma_{i}\succ\sigma_{j})\in\mathcal{P}$ that $\pi$ passes but all other policies in $\Pi$ do not pass.
@@
Definition 3 (Hidden context group). A hidden context group is a group of $m$ annotators, each with their own reward function $r_{1},r_{2},\dots,r_{m}$ that all monotonically rank the segments $\sigma\in\mathcal{D}$ the same.
@@
Definition 4 (Optimal policy for hidden context group). An optimal policy $\pi^{*}_{g}$ for a hidden context group $g$ is the optimal policy in a given environment (MDP/R) using the group’s implicit ground truth reward function $r_{g}$ as the reward function.
@@
Definition 5 (Contradictory preferences). A pair of preferences $(\sigma_{i}\succ\sigma_{j})$, $(\sigma_{x}\succ\sigma_{y})$ are contradictory under a specific policy regularization scheme if the likelihood of any policy that satisfies both is lower than the likelihood of a policy that satisfies either.
@@
With no regularization, a policy that satisfies two preferences should have a higher likelihood than one that satisfies one but not the other. This is because the likelihood of a certain policy is related to the total number of preferences passed by it. Passing two preferences would therefore illicit a higher likelihood than passing just one of them. However, if by satisfying both preferences, the models would need to incur a much greater regularization loss, it is likely that these two preferences came from individuals with differing hidden context.
@@
Theorem 1. In a completely noiseless setting, all policies that are optimal for specific HC groups are Pareto optimal with respect to the set of all preferences $\mathcal{P}$ generated from all the groups, and the space of all possible policies $\Pi=\{\pi|\pi\in\mathcal{S}\times\mathcal{A}\mapsto[0,1]\}$
@@
5.2 Lexicase Selection

### v2

# Pareto Optimal Learning from Preferences with Hidden Context

@@ Abstract
Ensuring AI models align with human values is essential for their safety and functionality. Reinforcement learning from human feedback (RLHF) leverages human preferences to achieve this alignment. However, when preferences are sourced from diverse populations, point estimates of reward can result in suboptimal performance or be unfair to specific groups. We propose Pareto Optimal Preference Learning (POPL), which enables pluralistic alignment by framing discrepant group preferences as objectives with potential trade-offs, aiming for policies that are Pareto-optimal on the preference dataset. POPL utilizes lexicase selection, an iterative process that selects diverse and Pareto-optimal solutions. Our theoretical and empirical evaluations demonstrate that POPL surpasses baseline methods in learning sets of reward functions and policies, effectively catering to distinct groups without access to group numbers or membership labels. We verify the performance of POPL on a stateless preference learning setting, a Minigrid RL domain, Metaworld robotics benchmarks, as well as large language model (LLM) fine-tuning. We illustrate that POPL can also serve as a foundation for techniques optimizing specific notions of group fairness, ensuring safe and equitable AI model alignment.
@@ 1 Introduction
For both safety and functionality, it is critical for AI models to align with the values of human users and stakeholders. Recently, reinforcement learning from human feedback (RLHF) (Christiano et al., 2017) has emerged as an effective mechanism for model alignment, using preferences to capture human values. However, when preferences are sourced from large groups of potentially diverse people, methods that rely on point estimates of human values are bound to either be suboptimal for all groups or unfair to certain groups, both of which are problematic in their own ways.
@@
In this work, we build upon the notion of hidden context proposed by Siththaranjan et al. (2023) and focus on the problem of Reinforcement Learning from Human Feedback with Hidden Context (RLHF-HC). Hidden context refers to information that is unavailable to a preference learning system yet affects the preferences given. For example, a person’s dominant hand might determine on which side they would prefer a robotic assistant to hand them an object. Under this formulation, our goal is to build a set of policies that contains the optimal policy under the reward function for each group of people. In practice, we see two clear use cases of such a set of policies. First, they can be selected from at test time to find an optimal policy for a given user without in a few-shot manner. Second, this set can be used to measure and ensure fairness between groups. Minimizing risk with respect to this diverse distribution of policies ensures that no specific group is disregarded in the risk measurement–thus enhancing safety.
@@
Preferences with hidden context may be contradictory i.e., not mutually satisfiable by a well-regularized policy or reward function. So, we propose to frame these preferences as objectives with potential trade-offs between each other. With this re-framing, the optimal policy for each individual hidden context group would be Pareto-optimal (non-dominated) on the dataset of preferences. With this in mind, we propose Pareto Optimal Preference Learning (POPL), where we learn a set of reward functions or policies (directly) that are optimized towards being Pareto-optimal with respect to the set of preferences given by a potentially diverse set of human annotators. To do this, we use an iterative selection process known as lexicase selection (Spector, 2012), which has been shown under mild assumptions to select individuals that are both Pareto-optimal and diverse. An outline of our method can be found in Figure 1. Our contributions can be summarized as follows:
  • We extend the problem of Reinforcement Learning from Human Feedback with Hidden Context (RLHF-HC) introduced by Siththaranjan et al. (2023), addressing critical limitations in preference learning for sequential, time-based domains, as opposed to contextual bandits.
  • We derive theoretical results proving that optimal reward functions and policies for hidden context groups are inherently Pareto-Optimal with respect to the given preferences, establishing a rigorous mathematical basis for our approach.
  • We develop “Pareto-Optimal Preference Learning” (POPL), a framework leveraging lexicase selection to generate a set of Pareto-Optimal reward functions or policies. POPL ensures diverse, group-specific alignment with human preferences, enabling robust personalization and fairness.
  • We demonstrated POPL’s superiority over strong baselines in diverse settings, including:
    • Minigrid RL: Policy learning in grid-based decision making dsomains.
    • Metaworld: Balancing safety and speed in 3D robotics manipulation tasks.
    • LLM Jailbreaking Detection: Mitigating harmful outputs by aligning preferences for both helpfulness and harmlessness (without labels).
  • We showcased POPL’s ability to efficiently scale to high-dimensional tasks, such as those involving LLMs, while maintaining computational efficiency. POPL achieves robust results with pre-trained models, making it broadly applicable across domains requiring fairness, alignment and diversity.
@@
Figure 1: An outline of the proposed Pareto Optimal Preference Learning (POPL) framework. Given a set of pairwise preferences over trajectory segments from groups with potentially different ground truth reward functions, we infer a set of reward functions or policies that captures each group’s ground truth, without group membership labels. To do this, we frame reward inference as multi-objective optimization, where each preference forms a single objective, and find a set of Pareto-optimal reward functions or policies.
@@ 2 Related Work
Data used for RLHF systems often comes from multiple people, who are diverse in their preferences and values (Bobu et al., 2023; Peng et al., 2023; Biyik & Sadigh, 2018; Santurkar et al., 2023). This data, when considered in its aggregated form, can not be captured perfectly by a decision-making model that relies on a point estimate of utility (Casper et al., 2023). These models try to find a single reward function that is most likely, which is often not the optimal reward function for any one single person. When the groups are not perfectly balanced, the minority groups might be underrepresented in the inferred reward function (Siththaranjan et al., 2023; Feffer et al., 2023; Kirk et al., 2023; Myers et al., 2021) or simply treated as noise (Baumler et al., 2023). There have been attempts at explicitly modeling different people with different levels of expertise (Gordon et al., 2021; Daniels-Koch & Freedman, 2022; Gordon et al., 2022; Barnett et al., 2023), but these methods generally rely on concrete ways to distinguish between groups.
Dealing with this Diversity
@@
In the context of RL, Myers et al. (2021) outlines an approach to learning a multi-modal reward function from online interaction between a human expert and a preference learning system. Ramé et al. (2024) learn a set of reward models by optimizing for diversity amongst the outputs. While similar to our approach, we also aim to align our reward models with hidden context groups through optimizing for Pareto-optimality. For generative AI models and LLMs, there have been a variety of studies attempting to align large models with diverse human preferences. Chakraborty et al. (2024) and Siththaranjan et al. (2023) learn a mixture of preference distributions or a parameterized reward distribution, respectively. However, both these techniques operate under a contextual bandit setting which results in sub-optimal performance when used in the more general RL setting (discussed further in Section 3). Bradley et al. (2024) and Ding et al. (2024) leverage fine-tuning to improve the diversity of model responses for better alignment and creativity, which do not directly address the ambiguity and hidden context in human preferences. Jang et al. (2023) and Dai et al. (2023) elicit preferences specifically along different dimensions in order to cater custom reward functions for users in test time, and to be safe with respect to conflicting objectives, respectively. Whilst we also aim to cater reward functions in test time as well as optimize fairness between groups, we do not have access to labels regarding the context of the preferences generated. Finally, Rame et al. (2024) also generates a set of Pareto-optimal reward functions. However, in their setting, the system has access to ground truth reward functions for each group, and the Pareto-front is generated through weight interpolation between these functions.
Bayesian Reward Inference
@@
Bayesian Reward Extrapolation (B-REx) (Brown et al., 2020b) instead learns a distribution of reward models from pairwise human preferences. B-REx is then able to perform Bayesian inference using MCMC (MacKay, 1992) to sample from the posterior of reward functions. With this distribution, a practitioner can establish high confidence performance bounds that can be used to assess risk in evaluated policies as well as detect reward hacking behaviors. However, B-REx and other reward inference methods often rely on a faulty assumption that humans provide preferences in a Boltzmann-rational way.
@@ 3 Preliminaries
Reinforcement learning from human feedback considers human preferences over trajectories (or more generally, outputs of a model) in order to learn a reward model or policy that respects the preferences (Brown & Niekum, 2019; Rafailov et al., 2024; Hejna et al., 2024; Casper et al., 2023; Finn et al., 2016).
In order to learn meaningfully from human preferences, one must characterize how preferences are generated from some parameterized preferences model $P(\sigma_{i}\prec\sigma_{j})$. Usually, this preference model is based on the notion of Boltzmann-rationality, where humans generate preferences in accordance to the Bradley-Terry (BT) model (Bradley & Terry, 1952). The probability of pairwise preference $(\sigma_{i}\succ\sigma_{j})$ between two trajectories segments given some utility function $f(\sigma)$ can be written as
@@
where $\beta$ models the confidence in the preference labels. $\beta\rightarrow\infty$ signals that the preference provider is perfectly rational, and $\beta=0$ signals that preferences are random. The BT model is used in many fields, such as psychology (Baker et al., 2009; Goodman et al., 2011; Goodman & Stuhlmüller, 2013). However, this model does not perfectly capture the mechanisms driving these preferences (Ghosal et al., 2023; Jeon et al., 2020; Knox et al., 2022; Bobu et al., 2020; Lee et al., 2021).
Hidden Context
Siththaranjan et al. (2023) introduce the problem of preference learning with hidden context. This is the idea that preferences are generated not only based on the exponential utility (partial return or regret), but also on some latent hidden context variable $z$. This variable is not accessible to preference learning systems and poses a challenge as it is often the case that this variable results in breaking the assumption that preferences are generated Boltzmann-rationally.
Marginalized Distributional Preference Learning
In order to account for hidden context in the preferences learned, Siththaranjan et al. (2023) introduce Distributional Preference Learning, which relies on a single model of utility $u(s|z)$ to output a distribution of utility assignments $u(s)$ for each state $s\in\mathcal{S}$, marginalizing over the hidden context variable $z$. In other words, they are able to represent the marginalized probability $P(R|s)$ of a specific utility $R$ in a state $s$. Herein, we will refer to a model that outputs a distribution per state as a Marginalized Distributional Preference Learning (MDPL) system.
Due to the marginalization process inherent in these systems, the utility function is unable account for persistent annotator identity—the fact that hidden context transcends a single preference annotation. In a contextual bandit setting such as those often found for finetuning LLMs (Rafailov et al., 2024), this is not an issue, as determining that an output has high risk simply depends on the distribution of rewards attributed to that specific state. However, in sequential tasks, where there are relationships between preferences at different times, it is important to maintain full, coherent, reward functions or policies for each group.
An example of how using an MDPL can lead to fairness issues is outlined in Figure 2. There are two groups that have different utilities for three states A, B and C. MDPLs fail to differentiate between trajectories like AB, BC, and AC, which have distinct fairness profiles and utilities for different groups, as their representation marginalizes over group-specific utilities.
Figure 2: An example of a situation where using POPL is preferable to using a Marginalized Distributional Preference Learning (MDPL) system. Due to the fact that these systems marginalize over the hidden context $z$ for each state, MDPLs are unable to be sensitive to persistent annotator identity. MDPLs represent the distribution of utility values in a column-wise fashion, or maintain a distribution of utilities for each state, that is decoupled from that for other states. Therefore, the utility for both groups of the trajectory $AB$ is indistinguishable from that for $BC$ by an MDPL. POPL, on the other hand, represents the distribution row-wise, finding a set of utility functions that should include the ground truth for each group. In this case, POPL can represent the fact that $AB$ is an unfair trajectory and $BC$ is fair, whereas MDPLs are unable to make this distinction.
@@
Contrastive Preference Learning (CPL) (Hejna et al., 2024) learns a policy directly from preferences without needing to learn an intermediate reward function. This method uses a regret-based model of preferences rather than the standard partial return interpretation. The probability of a preference under a candidate policy can be written as the ratio of the exponentiated sum of log-likelihoods of the chosen segment to the disregarded segment. We choose CPL over Direct Preference Optimization (DPO) Rafailov et al. (2024) as DPO can be derived as a special case of CPL with trajectories of length 1, starting from the same state. Furthermore, POPL is designed to be used in a variety of sequential, time-based domanis, but DPO and other contemporary RLHF methods restrict themselves to contextual bandit settings (such as in large language models). CPL, on the other hand, overcomes these limitations Hejna et al. (2024).
@@ 4 Problem Statement and Theoretical Foundation
In light of our discussion in section 3, we re-frame the problem of preference learning with hidden context as follows. The goal is to learn a set $\Pi=\{\pi_{1},\pi_{2}\cdots,\pi_{n}\}$ of policies such that, for the hidden context group represented by a variable $z\in\mathcal{Z}$, there is a policy $\pi_{z}\in\Pi$ that is the optimal policy for the ground truth reward function $r_{z}$ for the group. Note that this can be accomplished by standard (reward-based) RLHF (experiments in sections  6.1 and 6.4) or direct (reward-free) RLHF (experiments in sections 6.2 and 6.3). For the standard approach, a series of reward functions $R=\{r_{1},r_{2}\cdots,r_{n}\}$ are first learned from preferences, then used to train $n$ optimal policies. In the direct approach, the policies are learned directly from preferences such as done by Hejna et al. (2024) and Rafailov et al. (2024).
We now show that optimal policies for hidden context groups are Pareto-optimal with respect to the set of preferences given by all annotators. Therefore, recovering the set of pareto-optimal policies is a viable way to solve the RLHF-HC problem formatted above.
Definition 1 (Policy passing preference).
@@
A policy $\pi(a|s):\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ passes a preference $(\sigma_{i}{\prec}\sigma_{j})$ if the probability of the preferred segment $\prod_{(s,a)\in\sigma_{j}}\pi(a|s)$ is higher than the probability of the other segment $\prod_{(s,a)\in\sigma_{i}}\pi(a|s)$. Or, equivalently, if $\sum_{(s,a)\in\sigma_{j}}\log\pi(s,a)>\sum_{(s,a)\in\sigma_{i}}\log\pi(s,a)$.
Definition 2 (Policy-set-relative Pareto-optimality).
@@
A policy $\pi(a|s):\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ is Pareto optimal with respect to a set of preferences $\mathcal{P}$ relative to a set of other policies $\Pi=\{\pi_{1},\pi_{2},\cdots,\pi_{n}\}$ if and only if there exists a preference $(\sigma_{i}\succ\sigma_{j})\in\mathcal{P}$ that only $\pi$ passes out of all policies in $\Pi$
Definition 3 (Hidden context group).
@@
A hidden context group is a group of $m$ annotators, each with their own reward function $r_{1},r_{2},\dots,r_{m}$ that identically rank the segments $\sigma\in\mathcal{D}$.
Definition 4 (Optimal policy for hidden context group).
@@
An optimal policy $\pi^{*}_{g}$ for a hidden context group $g$ is an optimal policy in a given environment (MDP/R) using the group’s implicit ground truth reward function $r_{g}$ as the reward function.
Definition 5 (Contradictory preferences).
@@
A pair of preferences $(\sigma_{i}\succ\sigma_{j})$, $(\sigma_{x}\succ\sigma_{y})$ are contradictory under a specific policy regularization scheme if the likelihood of any policy that satisfies both is lower than the likelihood of a policy that satisfies either.
@@
With no regularization, a policy that satisfies two preferences should have a higher likelihood than one that satisfies one but not the other. This is because the likelihood of a certain policy is related to the total number of preferences passed by it. Passing two preferences would therefore elicit a higher likelihood than passing just one of them. However, if by satisfying both preferences, the models would need to incur a much greater regularization loss, it is likely that these two preferences came from individuals with differing hidden contexts.
Theorem 1.
@@
In a completely noiseless setting, all policies that are optimal for specific HC groups are Pareto optimal with respect to the set of all preferences $\mathcal{P}$ generated from all the groups, and the space of all possible policies $\Pi=\{\pi|\pi\in\mathcal{S}\times\mathcal{A}\mapsto[0,1]\}$
@@
5 Pareto Optimal Preference Learning
In this section, we outline an algorithm that can be used to generate a set of polices or reward functions that align with the preferences of different groups of people. We introduce lexicase selection, a method that can select candidate hypotheses that lie on the Pareto front. Our population represents a belief distribution that is updated based on observed evidence. Lexicase selection continually narrows the hypothesis space based on selection criteria, effectively ‘learning’ which policies or reward functions hold promise given the current (hidden-context-laden) data.
Lexicase Selection for Pluralistic Outcomes
Figure 3: Lexicase selection being used to select a single candidate hypothesis. Starting with a random ordering, the pool of reward hypotheses is filtered down based on the preferences in order, until a single individual remains or we run out of preferences. The resulting reward function is added to the next pool, and this process is repeated (with new shuffles) to fill the population.